# Compiler Optimization via LLM Reasoning for Efficient Model Serving

### v1

@@ Abstract
While model serving has unlocked unprecedented capabilities, the high cost of serving large-scale models continues to be a significant barrier to widespread accessibility and rapid innovation. Compiler optimizations have long driven substantial performance improvements, but existing compilers struggle with neural workloads due to the exponentially large and highly interdependent space of possible transformations. Although existing stochastic search techniques can be effective, they are often sample-inefficient and fail to leverage the structural context underlying compilation decisions. We set out to investigate the research question of whether reasoning with large language models (LLMs), without any retraining, can leverage the context-aware decision space of compiler optimization to significantly improve sample efficiency. To that end, we introduce a novel compilation framework (dubbed Reasoning Compiler) that formulates optimization as a sequential, context-aware decision process, guided by a large language model and structured Monte Carlo tree search (MCTS). The LLM acts as a proposal mechanism, suggesting hardware-aware transformations that reflect the current program state and accumulated performance feedback. Monte Carlo tree search (MCTS) incorporates the LLM-generated proposals to balance exploration and exploitation, facilitating structured, context-sensitive traversal of the expansive compiler optimization space. By achieving substantial speedups with markedly fewer samples than leading neural compilers, our approach demonstrates the potential of LLM-guided reasoning to transform the landscape of compiler optimization. 111Code is available at https://github.com/Anna-Bele/LLM˙MCTS˙Search**footnotetext: Equal contribution
@@ 1 Introduction
The rise of model serving for LLMs, diffusion models, and other neural models has enabled a new class of intelligent systems, driving transformative applications in healthcare, education, and scientific discovery. These models incur significant computational demands during inference, which proportionally translate into substantial monetary costs. Driving down the cost of model serving is critical, not merely to broaden access and democratize inference, but to catalyze faster cycles of innovation in model design and deployment. Achieving this goal demands reducing inference runtime on computational infrastructure, resources that are not only expensive but also increasingly limited in availability. Compiler optimizations are a critical enabler, not only for cost-efficient inferencing across diverse applications but also for empowering rapid research iteration. According to Bill Dally [1], NVIDIA’s Chief Scientist, compiler optimizations have driven over a 400 improvement in runtime over the past decade, compared to just 2.5 from semiconductor scaling.
@@
Existing compilers struggle with neural models due to the exponentially large space of valid program transformations (e.g., tiling, fusion, and layout changes). Each decision, such as selecting a tiling factor or a parallelization strategy, introduces dependencies and constraints that influence the feasibility and performance benefits of subsequent transformations. As a result, neural compilation has increasingly turned to stochastic search techniques, such as genetic algorithms and evolutionary strategies, for optimization [2, 3, 4, 5]. While these methods have shown promise in discovering performant configurations, they are fundamentally sample-inefficient. As a result, these techniques often explore redundant or invalid configurations. They also overlook synergistic transformations that emerge only when decisions are made with contextual awareness.
@@
We set out to investigate the research question of whether reasoning with large language models (LLMs), without any retraining, can leverage the context-aware decision space of compiler optimization to significantly improve sample efficiency. To that end, we introduce a novel compilation framework that couples large language model (LLM) reasoning with Monte Carlo tree search (MCTS) to guide compiler optimizations. Hence, in our approach, compiler optimization is cast as a sequential decision-making process, in which each transformation, such as tiling, fusion, or vectorization, is selected with awareness of the current program state, while assimilating downstream information and propagating its implications upstream to guide future decisions. Our approach avoids the prohibitive cost of fine-tuning LLMs as compilation policies and does not require additional training or task-specific adaptation. In this formulation, the LLM evaluates partial transformation sequences and proposes contextually appropriate next steps, drawing upon hardware-aware cost models and the historical trajectory of optimization decisions. The LLM serves as a context-aware proposal engine: given the current schedule and its observed performance, it generates candidate mutations that are likely to be effective in the context of the traversed trajectory. These LLM-guided reasoning choices are integrated into a Monte Carlo tree search (MCTS) framework, which provides a structured mechanism for balancing exploration and exploitation by evaluating LLM-suggested transformations, expanding promising branches, and leveraging rollout feedback to adaptively steer the search toward high-performing regions of the exponentially large optimization space.
@@
This integration of LLM-based chain-of-thought (CoT) guidance with tree search combines contextual reasoning and adaptability with principled, structured decision-making, enabling the compiler to navigate the complexity of the optimization search space with significantly improved sample efficiency. Empirical results show that our method achieves up to 2.5 speedup over unoptimized code using just 36 program samples, whereas state-of-the-art black-box autotuners like TVM using Evolutionary Search require up to 16 more program samples to reach comparable performance. These results underscore the promise of LLM-guided reasoning in neural compilation for efficient and scalable model serving.
@@ 2 Problem Formalization
To facilitate an efficient search over the space of valid program transformation sequences, we cast the optimization problem as a finite-horizon Markov decision process (MDP) defined by the tuple . This formulation provides a structured approach for sequential decision-making in the transformation space, allowing the search process to account for how individual transformations compound over time to affect final program performance. Compared to unstructured methods such as exhaustive or purely stochastic search, which often require a large number of expensive program evaluations, casting the problem as an MDP enables more deliberate exploration, offering the potential for improved sample efficiency. Each state corresponds to a program obtained by applying a sequence of transformations to the original program , i.e., . An action corresponds to selecting a transformation to apply at step , transitioning the current program to a new variant. Since the application of a transformation is deterministic, the transition function is if and 0 otherwise. The reward function is defined as the objective value caused by the optimization sequence, i.e., . By formulating the problem as an MDP, we enable the use of planning algorithms such as Monte Carlo tree search (MCTS) to explore program transformation sequences. The search ultimately yields an optimized transformation sequence that approximately maximizes the objective function, as defined in Equation (1).
@@ 3 Reasoning Compiler: Integrating LLM-Guided Contextual Reasoning with Monte Carlo Tree Search
We present the Reasoning Compiler, a novel compilation framework that unifies the structured exploration capabilities of Monte Carlo tree search (MCTS) with the contextual, history-aware reasoning of large language models (LLMs). While MCTS provides a principled approach to exploring sequences of program transformations, compiler optimization introduces a unique challenge: the successive application of transformations can exhibit complex, non-local interactions that are difficult to capture through purely stochastic or myopic policies. To address this, we employ an LLM to model program transformation context, tracking which transformations have been applied, how they impact performance, and what directions remain promising. This contextualization is essential to enabling effective and sample-efficient search in compiler optimization.
Optimization interaction is complex, complicating optimization. Unlike tasks where actions are relatively independent, program transformations compose in subtle and complex ways. For example, the profitability of applying loop tiling may depend on the prior application of loop fusion or unrolling. Additionally, transformations can introduce new, unforeseen opportunities/constraints for future transformations. These dependencies make the space of valid and useful transformation sequences both combinatorially large and deeply contextual. While black-box optimization methods such as evolutionary search and some implementations of reinforcement learning have achieved notable success in compiler autotuning [6, 2, 5, 7], they often do not explicitly model the nuanced structural and temporal dependencies between transformations. This can limit their ability to generalize across contexts where optimization efficacy depends on intricate transformation histories. Even when guided by local reward signals, they may struggle to capture the interplay between past decisions and future opportunities, limiting their effectiveness in deeply contextual optimization landscapes. Our insight is that efficient search in this space requires an agent capable of reasoning over the transformation history, structural code changes, and observed performance dynamics to inform the next transformation step.
@@ 3.1 LLM-Guided Contextual Reasoning for Program Transformation Proposal
Contextual reasoning via LLMs. To address these challenges, the Reasoning Compiler leverages a large language model (LLM) as a contextual reasoning engine. The LLM is tasked with synthesizing program transformation sequence proposals that are not only syntactically valid but also informed by the full history and structure of the program. By prompting the LLM with a rich, structured representation of the current optimization state, we enable it to reason over the cumulative effects of prior transformations, analyze performance trends, and identify differential improvements over prior programs.
@@
Figure 1 illustrates the optimization workflow. From the root program, the Reasoning Compiler traverses the tree by computing the UCT score [8], selecting a promising leaf node with corresponding program for expansion by balancing exploitation of high-reward paths and exploration of under-sampled branches based on visit statistics and node costs (see §3.2).
Prompt construction. At each expansion step in the search, the LLM receives a prompt that includes the source code and predicted performance cost for the current program , its parent , and its grandparent . It also includes the ordered sequences of transformations that were applied to reach each of these program variants, denoted , , and . Finally, the full set of available transformation operations is included. Given this context, the LLM is explicitly instructed to:
1.
Analyze the differences between program variants and their associated costs, identifying which transformations contributed to observed performance changes.
2.
Reason about potential interactions between previously applied and candidate future transformations, including both synergistic and antagonistic effects.
3.
Synthesize a new sequence of transformations that is justified in the context of the current program structure and transformation history.
4.
Provide a rationale for the proposed sequence, referencing specific code features, transformation interactions, and performance considerations.
This structured prompt is designed to elicit chain-of-thought (CoT) reasoning [9], encouraging the LLM to perform deep, multi-step analysis and move beyond surface-level edits, instead generating proposals that are both semantically meaningful and tailored to the evolving optimization trajectory.
@@
Transformation proposal and validation. The LLM proposes a candidate transformation sequence in the form of a string containing transformation names. Given the generative nature of the LLM, the output may include invalid or unrecognized transformations even though it is guided by a predefined set of valid transformations To ensure correctness, the output string is first parsed and filtered to retain only those transformations that match known valid names and transformation parameters. If no valid transformations remain after this filtering, we fall back to sampling a single random transformation from the valid set. The successfully validated and applied sequence yields a new program variant , with its transformation history updated as , where denotes sequence concatenation. This new program variant is scored using a hardware cost model and used to update the MCTS tree (see §3.2).
@@
It is important to emphasize that the LLM is not the centerpiece of our contribution, but a necessary enabler of effective search in this domain. Compiler optimization poses a uniquely challenging setting due to the non-local, compositional nature of transformation interactions. Traditional black-box search or heuristic-guided methods struggle to navigate such spaces efficiently. The Reasoning Compiler uses structured search (via MCTS) in conjunction with learned contextual reasoning (via LLM + CoT) to overcome these challenges. The result is a sample-efficient optimization algorithm capable of discovering performant transformation sequences in high-dimensional, high-interaction spaces.
@@ 3.2 Structured Optimization via Monte Carlo Tree Search
MCTS as a sample-efficient planner. As described in §2, we cast program optimization as a finite-horizon decision process over the space of transformation sequences. Framing the problem as an MDP allows the Reasoning Compiler to consider long-term transformation effects and invoke planning algorithms such as Monte Carlo tree search (MCTS) to explore this space deliberately and efficiently.
@@
MCTS operates over a tree where each node represents a program state corresponding to a program variant , and each edge corresponds to an action (i.e., a transformation sequence ). This tree structure naturally supports the reuse of common transformation prefixes and allows the planner to backpropagate value estimates from downstream program variants to upstream decisions. Such reuse is critical in compiler optimization, where transformation sequences exhibit both compounding effects and long-range interactions.
@@
Selection via UCT. During the selection phase, MCTS traverses from the root, recursively selecting children to maximize the UCT (Upper Confidence bounds applied to Trees) criterion:
@@
where is the estimated cost of , is the visit count of node , and governs the exploration-exploitation tradeoff.
@@
LLM-guided expansion. As shown in Figure 2(a), once a promising leaf node is selected, an LLM is queried to propose a transformation sequence conditioned on the program at and its ancestors (see §3.1). The model generates a candidate transformation sequence , which is applied to to produce a new program . This results in a new node added to corresponding to the updated program and extended transformation path. By leveraging the LLM’s contextual reasoning, the system proposes globally informed transformations that extend beyond myopic heuristics.
Simulation for local cost estimation. To assess the likely long-term consequences of the LLM-suggested transformation sequence at , the Reasoning Compiler simulates a short random trajectory starting from the new program . This is done by sampling a randomized sequence of legal transformations and applying them to obtain a terminal program . This simulates the potential downstream development of the current program state, modeling how the compiler might continue to optimize from onward under plausible future actions. The cost model then evaluates this trajectory to produce an estimated cost . This simulation step provides a noisy but informative proxy for the downstream impact of reaching , allowing MCTS to balance immediate and future consequences without expensive real-hardware runs.
As shown in Figure 2(b), once a new node is added to the tree, the Reasoning Compiler performs a lightweight MCTS simulation by randomly walking down a path to estimate the expected utility that further optimizing would attain in the future. This simulation is not a hardware simulation; rather, it is a common terminology used in MCTS algorithms. The performance of the program needs to be measured in these possible future nodes because the overall objective is to assess whether will eventually minimize runtime. Recall that the overall objective is to minimize a true hardware-level cost function (see §2). However, evaluating directly would require compiling and running the program on real hardware, which is an expensive and often impractical operation within the inner loop of a planning algorithm. To address this, the Reasoning Compiler approximates using a learned hardware-aware cost model , which is trained to mimic hardware performance while remaining significantly cheaper to evaluate. This surrogate model is denoted .
@@
Backpropagation. As illustrated in Figure 2(c), the estimated cost is then backpropagated to all ancestors along the path to the root according to the update step where is some ancestor. The visit counts are also updated according to the update step . These updates refine the empirical estimates that guide future selections.

### v3

# Reasoning Compiler: LLM-Guided Optimizations for Efficient Model Serving

@@ Abstract
While model serving has unlocked unprecedented capabilities, the high cost of serving large-scale models continues to be a significant barrier to widespread accessibility and rapid innovation. Compiler optimizations have long driven substantial performance improvements, but existing compilers struggle with neural workloads due to the exponentially large and highly interdependent space of possible transformations. Although existing stochastic search techniques can be effective, they are often sample-inefficient and fail to leverage the structural context underlying compilation decisions. We set out to investigate the research question of whether reasoning with large language models (LLMs), without any retraining, can leverage the context-aware decision space of compiler optimizations to significantly improve sample efficiency. To that end, we introduce a novel compilation framework (dubbed Reasoning Compiler) that formulates optimization as a sequential, context-aware decision process guided by a large language model and structured Monte Carlo tree search (MCTS). The LLM acts as a proposal mechanism, suggesting hardware-informed transformations that reflect the current program state and accumulated performance feedback. MCTS incorporates the LLM-generated proposals to balance exploration and exploitation, facilitating structured, context-sensitive traversal of the expansive compiler optimization space. By achieving substantial speedups with markedly fewer samples than leading neural compilers, our approach demonstrates the potential of LLM-guided reasoning to transform the landscape of compiler optimization.111Code is available at https://github.com/Anna-Bele/LLM˙MCTS˙Search**footnotetext: Equal contribution
@@ 1 Introduction
The rise of model serving for LLMs, diffusion models, and other neural models has enabled a new class of intelligent systems, driving transformative applications in healthcare, education, and scientific discovery. These models incur significant computational demands during inference, which proportionally translate into substantial monetary costs. Driving down the cost of model serving is critical, not merely to broaden access and democratize inference, but to catalyze faster cycles of innovation in model design and deployment. Achieving this goal demands reducing inference runtime on computational infrastructure, resources that are not only expensive but also increasingly limited in availability. Compiler optimizations are a critical enabler, not only for cost-efficient inferencing across diverse applications but also for empowering rapid research iteration.
@@
Existing compilers struggle with neural models due to the exponentially large space of valid program transformations (e.g., tiling, fusion, and layout changes). Each decision, such as selecting a tiling factor or a parallelization strategy, introduces dependencies and constraints that influence the feasibility and performance benefits of subsequent transformations. Rule-based optimizations also often rely on hand-tuned heuristics that can overfit to a specific workload or hardware target. The seminal work in superoptimization [superoptimization:asplos:1987, denali:pldi:2002, peephole-superoptimizer:asplos:2006] aimed to tackle these shortcomings through enumerative or symbolic search, but the search space proved combinatorial and rugged. STOKE [stoke:asplos:2013] showed that high-quality programs often lie in regions separated by low-probability paths, and therefore adopted Markov chain Monte Carlo (MCMC)-based randomized search. Neural compilers followed suit, using evolutionary search or simulated annealing to navigate similarly irregular landscapes [tensor-comprehensions:arxiv:2018, autotvm:neurips:2020, ansor:osdi:2020, metaschedule:neurips:2022]. While these methods have shown promise in discovering performant configurations, they are fundamentally sample-inefficient. They overlook synergistic transformations that emerge only when decisions are made with contextual awareness. These techniques also often explore redundant subspaces or invalid configurations.
@@
In contrast, we set out to investigate the research question of whether reasoning with large language models (LLMs), without any retraining, can leverage the context-aware decision space of compiler optimizations to significantly improve sample efficiency. To that end, we introduce a novel compilation framework that couples LLM reasoning with Monte Carlo tree search (MCTS) to guide compiler optimization. Hence, in our approach, compiler optimization is cast as a sequential decision-making process, in which each transformation, such as tiling, fusion, or vectorization, is selected with awareness of the current program state, while also assimilating downstream information and propagating its implications upstream to guide future decisions. Our approach avoids the prohibitive cost of fine-tuning LLMs as compilation policies, nor does it require additional training or task-specific adaptation. In this formulation, the LLM evaluates partial transformation sequences and proposes contextually appropriate next steps, drawing upon hardware-informed cost models and the historical trajectory of optimization decisions to inform its proposal. The LLM serves as a context-aware proposal engine: given the current schedule and its observed performance, it generates candidate transformations that are likely to be effective in the context of the traversed trajectory. These LLM-guided reasoning choices are integrated into an MCTS framework that provides a structured mechanism for balancing exploration and exploitation by evaluating LLM-suggested transformations, expanding promising branches, and leveraging rollout feedback to adaptively steer the search toward high-performing regions of the exponentially large optimization space.
@@
This integration of LLM-based chain-of-thought (CoT) guidance with tree search combines contextual reasoning and adaptability with principled, structured decision-making, enabling the compiler to navigate the complexity of the search space with significantly improved sample efficiency. We evaluate Reasoning Compiler and compare its improvements and sample efficiency with TVM, which employs evolutionary search. Results show that Reasoning Compiler consistently achieves significantly higher speedups than what TVM achieves using significantly fewer samples. On five representative benchmarks (Llama-3-8B Attention Layer, DeepSeek-R1 MoE Layer, FLUX Attention Layer, FLUX Convolution Layer, and Llama-4-Scout MLP Layer) and across five hardware platforms (Amazon Graviton2, AMD EPYC 7R13, Apple M2 Pro, Intel Core i9, and Intel Xeon E3), Reasoning Compiler achieves 5.0 average speedup using 5.8 fewer samples, resulting in an average of 10.8 improvement over TVM in sample efficiency. For the end-to-end Llama-3-8B benchmark across five hardware platforms, Reasoning Compiler uses 3.9 fewer samples to achieve a 4.0 speedup, yielding a 5.6 sample efficiency improvement. These results underscore the promise of LLM-guided reasoning in neural compilation for efficient and scalable model serving.
@@ 2 Problem Formalization
To facilitate an efficient search over the space of valid program transformation sequences, we cast the optimization problem as a finite-horizon Markov decision process (MDP) defined by the tuple . This formulation provides a structured approach for sequential decision-making in the transformation space, allowing the search process to account for how individual transformations compound over time to affect final program performance. Compared to unstructured methods such as exhaustive or purely stochastic search, which often require a large number of expensive program evaluations, casting the problem as an MDP enables more deliberate exploration, offering the potential for improved sample efficiency. Each state corresponds to a program obtained by applying a sequence of transformations to the original program , i.e., . An action corresponds to selecting a transformation to apply at step , transitioning the current program to a new variant. Since the application of a transformation is deterministic, the transition function is if and 0 otherwise. The reward function is defined as the objective value caused by the optimization sequence, i.e., .
By formulating the problem as an MDP, we enable the use of planning algorithms such as Monte Carlo tree search (MCTS) to explore program transformation sequences. Under standard assumptions, such as finite branching, bounded rewards, and a tree policy (e.g., UCT) that guarantees persistent exploration, MCTS is consistent on finite-horizon problems: as the number of simulations tends to infinity, it converges (with probability 1) to the optimal root action/sequence that maximizes the objective. With any finite simulation budget, it returns a high-quality but approximate solution. Consequently, our framework (see §3) yields a sequence that approximately maximizes the objective in practice while enjoying asymptotic optimality in theory.
@@ 3 Reasoning Compiler: Integrating LLM-Guided Contextual Reasoning with Monte Carlo Tree Search
We present Reasoning Compiler, a novel compilation framework that unifies the structured exploration capabilities of Monte Carlo tree search (MCTS) with the contextual, history-aware reasoning of large language models (LLMs). While MCTS provides a principled approach to exploring sequences of program transformations, compiler optimization introduces a unique challenge: the successive application of transformations can exhibit complex, non-local interactions that are difficult to capture through purely stochastic or myopic policies. To address this, we employ an LLM to model program transformation context, tracking which transformations have been applied, how they impact performance, and what directions remain promising. This contextualization is essential to enabling effective and sample-efficient search in compiler optimization.
Optimization interactions are complex, making efficient search challenging. Unlike tasks where actions are relatively independent, program transformations compose in subtle and complex ways. For example, the profitability of applying loop tiling may depend on the prior application of loop fusion or unrolling. Additionally, transformations can introduce new, unforeseen opportunities/constraints for future transformations. These dependencies make the space of valid and useful transformation sequences both combinatorial and deeply contextual. While black-box optimization methods such as evolutionary search and some implementations of reinforcement learning have achieved notable success in compiler autotuning [tvm, autotvm:neurips:2020, metaschedule:neurips:2022, chameleon:iclr:2020], they often do not explicitly model the nuanced structural and temporal dependencies between transformations. This can limit their ability to generalize across contexts, as optimization efficacy depends on transformation histories. Even when guided by local reward signals, they may struggle to capture the interplay between past decisions and future opportunities, limiting their effectiveness in deeply contextual optimization landscapes. Our insight is that efficient search in this space benefits from an agent that reasons over transformation history, structural code changes, and observed performance dynamics to choose the next step.
@@ 3.1 LLM-Guided Contextual Reasoning for Program Transformation Proposal
Contextual reasoning via LLMs. To address these challenges, Reasoning Compiler leverages a large language model (LLM) as a contextual reasoning engine. The LLM is tasked with synthesizing program transformation sequences that are not only syntactically valid but also informed by the full history and structure of the program. By prompting the LLM with a rich, structured representation of the current optimization state, we enable it to reason over the cumulative effects of prior transformations, analyze performance trends, and identify differential improvements over prior programs.
@@
Figure 1 illustrates the optimization workflow. From the root program, Reasoning Compiler traverses the tree by computing the UCT score [uct:ecml:2006], selecting a promising leaf node (i.e., program) for expansion by balancing exploitation of high-reward paths and exploration of under-sampled branches based on visit statistics and node costs (see §3.2).
Prompt construction. At each expansion step in the search, the LLM receives a prompt that includes the source code and predicted performance cost for the current program , its parent , and its grandparent . It also includes the ordered sequences of transformations that were applied to reach each of these program variants, denoted , , and . Finally, the full set of available transformation operations is included. Given this context, the LLM is explicitly instructed to: (1) analyze the differences between program variants and their associated costs, identifying which transformations contributed to observed performance changes; (2) reason about potential interactions between previously applied and candidate future transformations, including both synergistic and antagonistic effects; (3) synthesize a new sequence of transformations that is justified in the context of the current program structure and transformation history; and (4) provide a rationale for the proposed sequence, referencing specific code features and transformation interactions. This structured prompt is designed to elicit chain-of-thought (CoT) reasoning [cot:neurips:2022], encouraging the LLM to perform deep, multi-step analysis and move beyond surface-level edits, instead generating proposals that are both semantically meaningful and tailored to the evolving optimization trajectory.
@@
Transformation proposal and validation. The LLM proposes a candidate transformation in the form of a string. Given the generative nature of the LLM, the output may include an invalid or unrecognized transformation even though it is guided by a predefined set of valid transformations. To ensure correctness, the output string is first parsed and filtered to retain only a transformation that matches known valid names and transformation parameters. If no valid transformation is found, Reasoning Compiler samples a random transformation from the valid set. The successfully validated and applied transformation yields a new program variant , with its transformation history updated as , where denotes sequence concatenation. This new program variant is scored using a hardware cost model and used to update the MCTS tree (see §3.2).
@@
It is important to emphasize that the LLM is not the centerpiece of our contribution, but a necessary enabler of effective search in this domain. Compiler optimization poses a uniquely challenging setting due to the non-local, compositional nature of transformation interactions. Traditional black-box search or heuristic-guided methods struggle to navigate such spaces efficiently. Reasoning Compiler uses structured search (via MCTS) with learned contextual reasoning (via LLM + CoT) to overcome these challenges. The result is a sample-efficient optimization algorithm capable of discovering performant transformation sequences in high-dimensional, high-interaction spaces.
@@ 3.2 Structured Optimization via Monte Carlo Tree Search
MCTS as a sample-efficient planner. As described in §2, we cast program optimization as a finite-horizon decision process over the space of transformation sequences. Framing the problem as an MDP allows Reasoning Compiler to consider long-term optimization effects and leverage planning algorithms such as Monte Carlo tree search (MCTS) to explore this space deliberately and efficiently.
@@
MCTS operates over a tree where and such that each node is a program from the state space and each edge corresponds to a transformation from the action space . This tree structure naturally supports the reuse of common transformation prefixes and allows the planner to backpropagate value estimates from downstream program variants to upstream decisions. Such reuse is critical in compiler optimization, where transformation sequences exhibit both compounding effects and long-range interactions.
@@
Selection via UCT. During the selection phase, MCTS traverses from the root, recursively selecting child programs to maximize the UCT (Upper Confidence bounds applied to Trees) criterion:
@@
where is the estimated cost of , is the visit count of node (i.e., program) , and governs the exploration-exploitation tradeoff.
@@
LLM-guided expansion. As shown in Figure 2(a), once a promising leaf node is selected, an LLM is queried to propose a transformation conditioned on and its ancestors (see §3.1). The model generates a candidate transformation , which is applied to to produce a new program . This results in a new node added to corresponding to the updated program and extended transformation path. To ensure remains acyclic, if already exists in the tree, it is not added. By leveraging the LLM’s contextual reasoning, the system proposes globally informed transformations that extend beyond myopic heuristics.
Rollout for local cost estimation. As shown in Figure 2(b), once a new node is added to the tree, Reasoning Compiler performs a lightweight MCTS rollout to estimate the long-term impact of the transformation sequence that produced it. This is done by sampling a randomized sequence of legal transformations and applying them to obtain a terminal program . The objective is to minimize a hardware-level cost (see §2), but evaluating requires compiling and running on real hardware, which is too expensive for the inner loop of a planning algorithm. Following standard practice in compiler autotuning, Reasoning Compiler replaces with a learned, hardware-informed surrogate that is cheap to evaluate and has been shown to accelerate search while preserving final quality [tvm, autotvm:neurips:2020, ansor:osdi:2020, dynatune:iclr:2021, glimpse:dac:2022, transfer-tuning:pact:2022]. Therefore, the cost model evaluates this trajectory to produce an estimated cost . This noisy but informative proxy captures the downstream impact of reaching , enabling MCTS to balance immediate and future consequences without incurring real-hardware runs.
@@
Backpropagation. As shown in Figure 2(c), the estimated cost is then backpropagated to all ancestors along the path to the root according to the update step where is some ancestor program. The visit counts are also updated according to the update step . These updates refine the empirical estimates that guide future selections.